OpenStudy (anonymous):

Which of the following equations has only one solution? x2 + 4x + 4 = 0 x2 + x = 0 x2 - 1 = 0 and why

OpenStudy (tennistar):

They all do.... x2+4x+4=0 >> 6x = -4 >> x =-4/6 x2+x = 0 >> 3x =0 >> x =0 x2-1=0 >> 2x = 1 >> x =1/2

OpenStudy (anonymous):

when he says x2 he means x squared. and it has to be the first one, you can easily tell the other two have to solutions. But to make sure: \[x = { -4 \pm \sqrt{16-16} \over 2 } = -2\] So that is indeed the one.

OpenStudy (anonymous):

im a girl :P